Hello, my question is and i am having a lot of trouble with it even with 2 other friends is
A father is nine times the age of his son. in 5 years he will be 4 times the age of his son. how old are they now.
(i need the working out for it and it needs to be done using algebra help please!

Answer by josgarithmetic(39620)

"A father is nine times the age of his son. In 5 years, he will be 4 times the age of his son. How old are they now?

x = age of son now
y = age of father now

y=5x for "now".
-
Their ages five years from now are these:
x+5 = age of son in five more years
y+5 = age of father in five more years
y+5=4(x+5) ---- He will be four times the age of his son.

SUMMARY OF EQUATIONS:
----------------
y=5x
-
y+5=4(x+5)
----------------

SOLVING THE SIMULTANEOUS EQUATIONS
y+5=4x+20 from the 'in five years' equation;
4x+20=y+5
4x-y=5-20
y-4x=15
Substitute y using the 'now' equation;
5x-4x=15
1*x=15
x=15--------------part of the answer. The son, now.
-
Now compute y.
y=5x
y=5*15
y=75-------------other part of answer. The father, now.